OpenStudy (anonymous):

Which of the following is the equation of a line parallel to 4y=12x=5 that passes through (3,8)? A.y=3x-2 B.y+4=8x C.y+3x=2 D. y=4x-1

OpenStudy (anonymous):

1. let that line be \(y=mx+b\) 2. parallel lines have identical slopes. obtain \(m\) from the given line 3. this line passes throuogh (3,8) and hence, substitute \(x=3\) and \(y=8\) in the equation to find \(b\).

OpenStudy (anonymous):

So wait it would be 8=3m+b

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

first find "m"

OpenStudy (anonymous):

what is the slope of given line? (hint: ratio of y to x)

OpenStudy (anonymous):

Ive never been good at slopes.. So how would I find M

OpenStudy (anonymous):

original line: \(4y=12x=5\) now, \(m={y\over x}\)

OpenStudy (anonymous):

so it would be 8/3 fraction wise not dividing

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

\[4y=12x\\{y\over x}=? \]

OpenStudy (anonymous):

8 - 3

OpenStudy (anonymous):

\[12\over4\]!!!!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

do you see it?

OpenStudy (anonymous):

Oh!!!! I see now

OpenStudy (anonymous):

phew. so, \(m={12\over4}=3\)

OpenStudy (anonymous):

since our line is parallel to this line, its slope is the same good so far?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

nice so, with the work till now, we got \[y=3x+b\] so, we still need to find this "b" how? -> we use the point (3,8) plug in x=3 and y=8 in the above equation of line

OpenStudy (anonymous):

what value do you get for "b"?

OpenStudy (anonymous):

Ok so plugging it in 8=3(3)+b

OpenStudy (anonymous):

perfecto

OpenStudy (anonymous):

Okay so 8=9+b it would have to be a negative right

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

yes... it can be anything.. (positive or negative, dont matter)

OpenStudy (anonymous):

good. thats comples your answer \[y=3x-1\] what option says that?

OpenStudy (anonymous):

none

OpenStudy (anonymous):

I have one that says y=3x-2

OpenStudy (anonymous):

check the point and the original line eq.

OpenStudy (anonymous):

it says 4y=12x+5

OpenStudy (anonymous):

aha but that is fine.. (slope depends on 'y' and 'x'.. const term gives you the y-intercept) what about the point

OpenStudy (anonymous):

I don't get what you mean.. Everything that is on my paper I put up on the description

OpenStudy (anonymous):

math doesnt lie. so, check if you noted down the options and the point correctly.

OpenStudy (anonymous):

among the options, only the (A) line is parallel to the given line but it does not pass through (3,8)

OpenStudy (anonymous):

see if the point is (3,7)!!!!

OpenStudy (anonymous):

Ohhh I accidentally put 8 instead of 7 my bad

OpenStudy (anonymous):

see, told ya, math cannot lie, :)

OpenStudy (anonymous):

Haha. My bad

OpenStudy (anonymous):

So do we start all over or is it y=3x-2

OpenStudy (anonymous):

not.. just the "b" part

OpenStudy (anonymous):

plug in the (3,7) and find "b"

OpenStudy (anonymous):

ok so it would be 7=3m+b

OpenStudy (anonymous):

yes.. but you already have your "m" remember?

OpenStudy (anonymous):

Ok so m stays the same

OpenStudy (anonymous):

7=3(3)+b 7=9+-2 7=7 y=3x-2

OpenStudy (anonymous):

Am i right or anywhere close

OpenStudy (anonymous):

hold it.... you are "solving for b" \[7=3(3)+b\\\implies7-9=b\\\implies -2=b\]

OpenStudy (anonymous):

and now you have \[\boxed{y=3x-2}\]

OpenStudy (anonymous):

Ok can you help me with 2 more ones?